1. (2 points) calculate the [H30 ]for the following conditions a. pH 3.21 b. pOH 13.34 2. (4 points) calculate the pH of the following solutions: a. 3.7 x 10-3 M HCl b. 8.6 x 10-2 M HC H302 K 1.74 x 10 3. (2 points) Calculate the pH of the following buffer a. 7.8 x 10-2 M Hc2H,02 and 8.6x 10-2 M Nacz H302 4. (4 points) Abuffer is made from 3.7 x 10-1 M HCH,02 and 2.9 x 10 1 M NaC2H302 a. What is the pH? b. If you added 3.0 x 10-3 moles of Na0H to 0.100 Lof buffer what is the final pH. (assume that any change in volume is negligible.) 5. (2 points) write out the full ionization equations for the following amino acids. (Hint: use the last page of the pH review to help with the order of ionizations and start at neutral pH) a. Alanine b. Serine 6. (2 points) How many grams of phenylalanine are needed to make 350.0 ml of a 0.125M solution? 7. (4 points) You titrate valine as we did in class today; dissolved it, acidified the solution with an excess of HCI, then titrated with NaoH until a pH of 13.5 was reached. Use the following information to determine; a. The molecular weight of valine b. Calculate the pl Mass of Valine: 0.203g Molarity of NaOH 0.165M Not seen clearly Volume at equivalence point 1: pKa1 Volume at equivalence point 2: Pi 18.39 mL Volume at equivalence point 3: Ka2 28.91 ml 2.27 (estimated) Ka 1 Ka 2 9.68
Answer #1

1) a) pH = -log[H3O+]= 3.21

Thus, [H3O+] = 10-3.21 = 6.2 x 10-4 M

b) pOH = 13.34

we know that, pH +pOH = 14

Thus, pH= 14 - 13.34 = 0.66

-log[H3O+]= 0.66

Thus, [H3O+] = 10-0.66 = 0.22 M